how many terms of the sequence 13,11,9..... makes the sum 45? Explain the answer
Answer by swincher4391(1107)

sum from i =1 to n of (15-2i) = 45
sum from i = 1 to n of 15 - 2* sum from i = 1 to n of i
15n - n(n+1) = 45
15n - n^2 + n = 45
n^2 -14n +45 = 0
(n-5)(n-9)
n=5, n=9
Check: 13+11+9+7+5 = 45
Check: 13+11+9+7+5+3+1+-1+-3 = 45
So 5 terms or 9 terms.